A chemist has one solution that is 20% saline and a second that is 65% saline. How many gallons of each should be mixed to get 120 gallons of a solution which is 50% saline?
Answer by drk(1908)

This is a mixture problem. Here is the table:
SALINE . . . . . % . . . . . .gal . . . . . .%gal
A . .. . . . . . .. 20 . . . .. . G . . . . . . . 20G
B . . . . . . . . .. 65 . . . . .120-G . . . .7800 - 65G
mix . . . . . . . 50 . . . . . .120 . . . . . .6000.
--
I put in all information from the problem and then I had to use variables. I let G = gallons of A and 120-G be gallons of B. It is always total gallons - X.
Now,
20G + 7800 - 65G = 6000
-45G = -1800
G = 40.
I need 40 gallons of the 20% saline solution and 80 gallons of 65% saline solution.